OpenStudy (anonymous):

Vectors Given OA=2i-3j,OB=3i-6j and OC=5i-12j,show that the points A,B and C are collinear. @ganeshie8

OpenStudy (phi):

can you find the "direction vector" from OA to OB ?

OpenStudy (anonymous):

yes\[AB=\left(\begin{matrix}1 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

is it correct?

OpenStudy (phi):

the j component is -6 - (-3)

OpenStudy (anonymous):

ops\[AB=\left(\begin{matrix}1 \\ -3\end{matrix}\right)\]

OpenStudy (phi):

the equation of a line can be written as \[ p= p_0 + t \vec{v} \] where v is the direction vector, t is a scalar variable and p0 is a point on the line so the eq for a line with points A and B is \[ p= (2, -3) + t ( 1, -3) \] collinear means line on the same line. so we have to show that point (5, -12) is on the line. I would show there is a t that "works" i.e. solve for the i component 2+ t = 5 and show that t=3 works

OpenStudy (phi):

2i-3j + 3 * (i -3j) = 2i-3j + 3i -9j = 5i -12j

OpenStudy (phi):

all 3 points satisfy the equation (with t=0, t=1, and t=3 )

OpenStudy (anonymous):

can u tell me how to get t=0 and t=1 @phi

OpenStudy (phi):

when you find the equation P0 + t * v if we use point A as P0, (which is what we did), then t=0 gives us P0, i.e. gives us A but even if you did not know this, you would write (2, -3) + t (1, -3) = (2, -3) now pick either dimension (or component) for example, if we look at the "j" component (2nd dimension) we have this equation -3 + t*-3 = -3 add +3 to both sides t*-3= 0 t= 0 you can do that same thing for the 2nd point (3, -6) to find t= 1 works but because we used points A and B to find this equation it always turns out t=0 and t=1 will work to "get back" A and B

OpenStudy (anonymous):

Thank you @phi :)

OpenStudy (phi):

of course, once we find t, we have to verify that t gives us back the entire point. for example if we had (2, -3) + t (1, -3) = (2, -4) and we solve using i component: 2 + t*1 = 2 we find t=0 now we use t=0 in the equation (2,-3) + 0* (1,-3) = (2+0, -3+0) or just (2, -3) and that does not equal (2,-4) this shows there is no t that works for all of the components at the same time and that means that the point (2, -4) is not on the line

OpenStudy (baru):

We could also show all 3 vectors make the same angle with the x axis by using dot product

OpenStudy (phi):

***(2, -3) + t (1, -3) = (2, -4)*** that is an example to show a point that is not on the line. I picked (2,-4) as an arbitrary point that I know is not on the line.

OpenStudy (anonymous):

i c

OpenStudy (anonymous):

The book answer show that \[AB=\frac{ 1 }{ 3 }AC\],is it acceptable if i put \[AB=\frac{ 1 }{ 2 }BC\]

OpenStudy (phi):

yes

OpenStudy (anonymous):

Okay,thnx @phi :)